### v1

# Benchmark for Antibody Binding Affinity Maturation and Design

Abstract
We introduce AbBiBench (Antibody Binding Benchmarking), a benchmarking framework for antibody binding affinity maturation and design. Unlike existing antibody evaluation strategies that rely on antibody alone and its similarity to natural ones (e.g., amino acid recovery rate, structural RMSD), AbBiBench considers an antibody–antigen (Ab-Ag) complex as a functional unit and evaluates the potential of an antibody design binding to given antigen by measuring protein model’s likelihood on the Ab-Ag complex. We first curate, standardize, and share 9 datasets containing 9 antigens (involving influenza, anti-lysozyme, HER2, VEGF, integrin, and SARS-CoV-2) and 155,853 heavy chain mutated antibodies. Using these datasets, we systematically compare 14 protein models including masked language models, autoregressive language models, inverse folding models, diffusion-based generative models, and geometric graph models. The correlation between model likelihood and experimental affinity values is used to evaluate model performance. Additionally, in a case study to increase binding affinity of antibody F045-092 to antigen influenza H1N1, we evaluate the generative power of the top-performing models by sampling a set of new antibodies binding to the antigen and ranking them based on structural integrity and biophysical properties of the Ab-Ag complex. As a result, structure-conditioned inverse folding models outperform others in both affinity correlation and generation tasks. Overall, AbBiBench provides a unified, biologically grounded evaluation framework to facilitate the development of more effective, function-aware antibody design models.
1 Introduction
Antibodies are critical components of the adaptive immune system, functioning primarily by recognizing and binding specifically to antigens such as pathogens or aberrant cells. This specific recognition is facilitated through complementary regions: the antibody’s paratope and the antigen’s epitope. Enhancing the binding affinity between an antibody and its corresponding antigen can significantly increase therapeutic potency and diagnostic sensitivity, making antibody design and optimization a significant area of drug discovery. Developing a therapeutic monoclonal antibody depends on multiple factors, including expression, stability, immunogenicity, aggregation, and binding affinity [9, 30]. Among these, binding affinity is the most critical determinant of therapeutic efficacy, as it directly influences antibody potency. Insufficient binding affinity fundamentally limits an antibody’s utility. Therefore, increasing binding affinity between antibody and antigen has been a crucial process in therapeutic antibody development.
Traditional antibody discovery methods, such as phage display technology [48, 50, 65] and animal immunization [21, 37], employ iterative cycles of mutation and selection to progressively improve binding affinity. These methods have significantly advanced therapeutic antibody discovery but face challenges due to the initially limited diversity of libraries in phage display or the narrow naive B cell repertoire available in animal immunization models, restricting the comprehensive exploration of potential high-affinity antibody variants (Fig. 1). Machine learning-based antibody design approaches complement these experimental techniques by efficiently navigating the vast search space (Fig. 1) and proposing high-affinity antibody variants that may not be readily accessible through experimental approaches alone.
Figure 1: Antibody design space. (a) Traditional in vitro screening explores a limited antibody library. (b) Protein machine learning models explore broader mutational space. $\bullet$: tested antibodies, $\bullet$: lead antibody, $\circ$: model-generated variants.
Recent advances in machine learning, especially with protein language models and structure-based generative models [12, 24, 26, 27, 38, 39, 41, 45, 47, 49, 61, 63, 67, 75, 78], have shown promise in antibody design. However, common evaluation metrics like amino acid recovery rates or structural RMSD to natural antibodies do not adequately capture biological relevance in antibodies. In general protein design, comparing to the closest natural variant is a reasonable validation strategy because protein mutations are driven by strong evolutionary pressure. In contrast, antibodies are uniquely generated through semi-random process such as V(D)J recombination and somatic hypermutation, resulting in extraordinarily high sequence diversity. It is common for two antibodies that bind the same antigen to share minimal sequence similarity, and exceedingly rare that two arbitrary antibodies are similar unless they are from the same lineage of a human immune response. As such, evaluating designed antibodies by how closely they resemble naturally occurring ones overlooks the fundamental biology of antibody generation. This calls for new evaluation criteria that better reflect the functional goals of antibody design, rather than assumptions borrowed from general protein modeling.
From a structural biology perspective, binding affinity is determined not just by the antibody sequence, but by the quality of the interface it forms with the antigen. High-affinity binding typically arises from antibody-antigen (Ab-Ag) complexes that exhibit structural integrity [63] – meaning they are stable, well-packed, and maintain favorable conformations with minimal strain. Structural integrity ensures optimal shape and chemical complementarity at the binding interface. Antibodies that form such stable complexes with their targets resemble naturally occurring Ab-Ag complexes collected in structural database SAbDab [15]. Recent machine learning models [12, 27, 41, 45, 47, 67, 78] have shown success in learning Ab-Ag complex sequence-structure patterns, enabling us to gauge whether a designed Ab-Ag complex lies within the high-probability manifold of structurally stable, high-affinity complexes. Therefore, incorporating the antigen into evaluation provides a more biologically grounded and functionally relevant assessment of antibody design.
To address above limitations, we introduce AbBiBench (Antibody Binding Benchmarking), a biologically relevant benchmarking framework specifically designed for antibody binding affinity maturation. Rather than assessing antibodies in isolation [9], we consider the Ab-Ag complex as a unit for evaluation. We curated standardized data from publicly available experimental binding affinity studies, compiling 155,853 mutated heavy chain antibodies across nine antigen targets to evaluate protein models for binding affinity optimization. [^1] https://huggingface.co/datasets/AbBibench/Antibody_Binding_Benchmark_Dataset We also devised and publicly shared an efficient pipeline to rank newly designed antibodies based on complex structural integrity and biophysics properties. [^2] https://github.com/MSBMI-SAFE/AbBiBench AbBiBench is curated to avoid data leakage: although wild-type antibodies or antigens may appear in public datasets, no training corpus contains the mutant antibody–antigen complexes it evaluates. By providing a rigorous, biologically grounded benchmark for antibody design, AbBiBench will accelerate methods that lead to clinically and diagnostically impactful discoveries.
Figure 2: Overview of AbBiBench benchmarks. Antibody variants with experimentally determined affinity values are curated. Data modalities include amino acid sequences, wild-type antibody-antigen complexes, and affinity scores. A diverse set of baseline models includes general protein language models and specialized antibody models. All models are evaluated on two tasks: affinity prediction and antibody redesign. Five computational metrics assess newly designed antibodies from sequence plausibility, structural integrity, and binding affinity perspectives.
2 Related Work
2.1 Measuring Binding Affinity Between Antibody and Antigen
Three main approaches exist: experimental [1, 33, 43], biophysics-based [6, 8, 76], and data-driven (machine learning) methods [12, 18, 34, 42]. The most direct method to measure binding affinity is through the dissociation constant (Kd), which experimentally quantifies how tightly an antibody binds to its antigen. They are costly and low-throughput, thus computational models aim to approximate the binding affinity to enable large-scale screening. Biophysics-based computational models estimate Ab-Ag binding affinity by calculating interaction energies and free energy change, such as FoldX and $\Delta\Delta G$. It provides mechanistic insights into binding affinity. Though not direct affinity estimator, structural integrity (or plausibility) of Ab-Ag complex have been used as a proxy for binding affinity, measured via structure-based confidence metrics (interface pLDDT, iPTM) [18, 34] or ML-based perplexity (e.g., perplexity from AntiBERTy, ProteinMPNN, or ESM-IF) [12, 42, 61]. Despite the efforts to measure binding affinity computationally, the reliability and accuracy of these metrics remain an ongoing challenge [32].
2.2 Prior Benchmark Studies Evaluating Binding Affinity via Model Likelihood
Alternatively, some studies retrospectively evaluate machine learning models using historical experimental protein fitness data [9, 70, 72]. These works assume that higher model likelihoods correlate with higher protein fitness, suggesting the model can generate functional sequences. ProteinGym [55] exemplifies this approach by benchmarking models against experimental data on enzymatic activity, binding, expression, and stability. It measures correlation between model perplexity and fitness scores. However, ProteinGym does not include antibodies. FLAb [9] addresses this gap by compiling experimental measurements specifically for antibodies, including binding affinity, specificity (e.g., polyreactivity), immunogenicity, and developability metrics. While comprehensive, FLAb does not consider the antigen when evaluating binding affinity. This is a critical limitation, as antibody-antigen interactions are highly specific, and accurate affinity prediction requires modeling the Ab-Ag complex.
3 AbBiBench: An Antibody Binding Benchmark
We introduce AbBiBench, a benchmark to evaluate protein models’ ability to predict and design high-affinity antibodies. We assess zero-shot correlation between model likelihoods and experimental binding affinities (Sec. 3.2) across curated datasets (Sec. 3.1), and validate generative performance by designing CDR-H3 variants that improve F045-092 binding to H1N1 influenza (Sec. 3.3.2).
3.1 Datasets
We compiled 11 datasets of antigen, antibody heavy chain sequences, antibody structure, and experimental binding affinity measurements (Table 1). We focus on the mutation in variable heavy chain (vH) due to its high diversity from V(D)J recombination and the central role of CDR-H3 in antigen binding [7]. The vH often acts as a unique antigen-specific signature [13], while the light chain remains relatively conserved across functional antibodies [29], making vH the most relevant region for affinity optimization. For antibody structure, we focus on variable regions of heavy and light chain, where affinity conferring mutations occurs. We only included datasets with at least 20 mutated antibodies to ensure the statistical significance of evaluation.
In each binding affinity study, antibody libraries are constructed through phage or yeast display, introducing mutations via deep mutational scanning or at targeted positions. Some libraries were designed computationally using machine learning [40, 62] or biophysical modeling [10], resulting in variant sets ranging from 40 to 65,535. For consistency, we standardized experimental affinity measurements (Kd and enrichment ratio) by taking the negative log (for Kd) or log (for enrichment), so that higher values indicate stronger binding (Supplement LABEL:supplementary:dataset, Table LABEL:tab:dataset_details). While log enrichment is an indirect measure, log enrichment reflects how well a variant is retained after antigen-specific selection and correlates with binding strength. When normalized, it provides a scalable proxy for relative binding affinity in high-throughput screens. Notably, AbBiBench is designed to avoid data leakage: while wild-type antibodies and antigens may occur in public datasets such as OAS [57] and SAbDab [15], the specific antibody mutant–antigen complexes curated in our benchmark are not present in any known training corpus.
PDB ID Antibody Antigen Heavy-chain variants Binding score Study 4fqi_h1 CR9114 Influenza A/ New Caledonia/20/99 (H1N1) $65,094$ $-\!\log K_{d}$ [60] 4fqi_h3 CR9114 Influenza A/ Wisconsin/67/2005 (H3N2) $65,535$ $-\!\log K_{d}$ [60] 3gbn_h1 CR6261 Influenza A/ New Caledonia/20/99 (H1N1) $1,887$ $-\!\log K_{d}$ [60] 3gbn_h9 CR6261 Influenza A/ Hong Kong/1073/1999 (H9N2) $1,842$ $-\!\log K_{d}$ [60] aayl49 AAYL49 Spike HR2 $4,312$ $-\!\log K_{d}$ [17] aayl49_ML AAYL49_ML Spike HR2 $8,953$ $-\!\log K_{d}$ [40] aayl51 AAYL51 Spike HR2 $4,320$ $-\!\log K_{d}$ [17] 2fjg G6.31 VEGF $2,223$ log enrichment [36] 1mlc D44.1 Hen-egg-white lysozyme $1,297$ log enrichment [74] 1n8z trastuzumab HER2 $419$ $-\!\log K_{d}$ [62] 1mhp AQC2 Integrin-$\alpha$-1 $40$ $-\!\log K_{d}$ [10]
Table 1: Overview of the 11 Ab-Ag binding-affinity assays reported in AbBiBench, showing the number of heavy-chain mutants in each study and the respective binding metric.
3.2 Protein models
Protein modeling is a fast-evolving active research area. We selected diverse pretrained protein and/or antibody models based on originality, code availability, and structure modality (Table LABEL:tab:protein_model_summary).
Masked Language Models: Masked protein language models (MLMs) predict masked residues based on context and capture correlations between sequence motifs and higher-level functional properties, enabling their application in antibody design [25, 51]. We evaluate several representative protein MLMs. ESM-2 [42] is trained on large-scale protein sequence datasets using a masked language modeling objective. AntiBERTy [61] is a 12-layer BERT model trained on 57 million heavy- and light-chain sequences from antibody database (OAS [57]). Incorporating structural information into protein language models (PLMs) improves their ability to capture spatial context beyond sequence proximity. SaProt [67] extends ESM-2 with structure-aware tokens from Foldseek [4, 71], embedding residue identity and local structure. ProSST [41] uses geometric vector perceptrons (GVP) encoder [31] that discretizes local atomic neighborhoods into a compact codebook, with disentangled attention over sequence, structure, and position. ESM-3 [24], an upgraded ESM-2, is a multimodal PLM that models sequence, structure, and function through discrete token tracks. Based on such general-purpose protein MLMs, several antibody-specific MLMs have been developed, including CurrAb [5], AbLang [58], and IgBlend [47], to capture antibody-specific mutations driven by somatic recombination and hypermutation [61]. CurrAb, a fine-tuned ESM-2, uses curriculum learning to gradually shift from unpaired to paired OAS [57] antibody data while preserving pre-trained knowledge on general proteins.
Autoregressive Protein Language Models: Unlike MLMs, which predict masked tokens based on bidirectional context, autoregressive PLMs generate the next token using only left-to-right context. ProGen2 [54] is a Transformer-decoder model that scales up to 6.4 B parameters and is trained on $\sim\!1$ billion natural protein sequences. ProtGPT2 [19] adopts the GPT-2 architecture with 738 M parameters and is trained end-to-end on $\sim\!50$ million protein sequences. Inverse folding models are also autoregressive but based on global structure embeddings. Therefore, ProGen2 and ProtGPT2 can serve as structure-agnostic comparisons for inverse folding models.
Inverse Folding Models: Inverse folding models aim to predict amino acid sequences from a given protein structure, often in an autoregressive manner. This approach has been applied to antibody mutant design by leveraging known Ab-Ag complex structures [63], enabling sequence exploration to identify mutations that preserve or enhance complex stability and binding affinity. Widely used models include ProteinMPNN [12] and ESM-IF [27]. ProteinMPNN uses a message-passing neural network [20] to model residue interactions before autoregressive sequence generation. ESM-IF1 combines a GVP encoder for extracting backbone-invariant features with a Transformer decoder. AntiFold [26], based on ESM-IF1, is fine-tuned on antibody structures from SAbDab and OAS.
Diffusion-Based Generative Models: Diffusion models approach antibody design as a denoising process, transforming Gaussian noise into a target antibody by learning the joint distribution of atomic coordinates, orientations, and residue identities. DiffAb [45] conditions on an antigen–antibody framework complex and jointly diffuses CDR sequence and structure. To assess backbone flexibility, we also evaluated fixed-backbone variants DiffAb_fixbb, a common setting in protein design [3, 27, 28, 45, 66, 69]. AbDiffuser [49] extends to full-atom generation with physical priors for side chains. IgDiff [11] performs de novo backbone generation by sampling variable-region backbones and then fills in the sequences using AbMPNN [14].
CDR Imputation in Geometric Representation: Geometry-aware methods view affinity maturation as filling in missing CDRs within the explicit 3-D Ab-Ag interface. MEAN [38] masks CDRs on an E(3)-equivariant residue–atom graph containing both chains and epitope; two alternating message-passing blocks jointly restore CDR sequence and backbone. dyMEAN [39] upgrades this to full-atom, end-to-end design: conserved-framework initialization plus a “shadow paratope” lets the network emit paratope sequence, side-chain geometry, and binding pose in one shot. We also considered fixed-backbone versions of MEAN and dyMEAN, dubbed MEAN_fixbb and dyMEAN_fixbb, respectively.
3.3 Evaluation Tasks
Our benchmark comprises (i) zero-shot affinity prediction using retrospective experimental affinity data and (ii) antibody generation by sampling from the models.
3.3.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log-Likelihood
To measure how well a model’s zero-shot predicted log-likelihood aligns with wet-lab verified affinity, we calculated the Spearman correlation between the model likelihood and experimentally measured binding affinity. A high correlation indicates that the model assigns a higher likelihood to strong binders, suggesting that it can identify affinity-enhancing mutations in a zero-shot setting. To evaluate how effectively a model prioritizes the most promising antibodies, we also reported 5-fold precision@10 – the proportion of top 10 ranked variants that achieve at least 5-fold improvement in binding affinity compared to the wild type. The calculation of affinity fold change is detailed in Supplement LABEL:sup:sec:computationalMetrics. We harmonized the log-likelihood computation for all models under a unified setting: the input unit is the mutant-antigen complex, and the output is the likelihood of that complex. The details of likelihood computation across different types of models are provided in Supplement LABEL:calculation_details. In addition to the model’s log likelihood, we report two biophysics-based affinity score as a baseline: binding free energy ($\Delta G$) and the relative solvent-accessible surface areas ($SASA$) of epitope residues (Sec. 3.3.2, Supplement LABEL:sup:sec:computationalMetrics). Lower values of both metrics imply stronger binding. To ensure consistent directionality with model log-likelihoods, we report $-\Delta G$ and $-SASA$.
3.3.2 Generate Antibody Variants with Strong Affinity to H1N1 Influenza Virus
To assess whether protein models can generate antibody variants with improved binding to a specific antigen, we conducted a case study using F045-092 [56], a naturally occurring antibody that targets the hemagglutinin (HA) protein of influenza H3N2. Notably, F045-092 fails to bind the H1N1 strain California2009 due to a steric barrier at the HA receptor-binding site [16, 64]. No experimentally determined structure of the F045-H1N1 complex exists in public databases, and this particular pair has not been studied in prior antibody design literature. As such, our study represent a completely novel setting, free from data leakage. The structure used in our study was computationally predicted using AlphaFold3. We used four representative models—ESM-IF, SaProt, DiffAb, and MEAN—to redesign the CDR-H3 loop of F045-092 while allowing up to five substitution to maintain H3N2 cross-reactivity. Each model generated 1,500 CDR-H3 variants from an input consisting of the masked F045-092 sequence and a predicted complex structure with the H1N1 HA1 protein (Supplement LABEL:sup:sec:case_study). Sampling strategies varied by model, including autoregressive prediction (SaProt), greedy heuristics (ESM-IF), and diffusion-based or graph-based generation (DiffAb, MEAN; see Supplement LABEL:sup:sec:sec:design_cdrh3 for details).
Figure 3: Evaluation of new antibody design. Left: assess sequence plausibility and binding energy change. Right: high-confidence ranking based on AF3-predicted complex structures, evaluating complex structure confidence, epitope SASA, and inverse folding likelihood.
We evaluated the generated antibody variants from two perspectives: sequence plausibility and binding potential. Sequence plausibility was assessed using the log-likelihood from AntiBERTy [61], reflecting how closely mutations align with natural antibody evolution. We also computed inverse folding likelihoods using ProteinMPNN to determine whether each sequence is compatible with its backbone structure—higher scores indicate greater foldability. Binding potential was evaluated using three structure-based metrics: (1) binding free energy ($\Delta G$) to estimate Ab–Ag interaction strength, (2) epitope SASA to measure differences in the solvent-accessible antigen surface area upon binding, and (3) complex structure confidence (AlphaFold pLDDT [2]) as a proxy for interface stability (Supplement LABEL:sup:sec:computationalMetrics).
Some metrics require the structure of the antibody variant in complex with antigen. Predicting the complex structure of all sampled variants using a folding model is computationally expensive. To address this, we used a two-phased screening approach (Fig. 3). In Phase 1 we evaluated all 1,500 variants per model using sequence plausibility (AntiBERTy likelihood) and $\Delta G$ – metrics that do not require structure prediction – and selected top 20% antibody variants. In Phase 2, we generated full Ab-Ag complex structures with AlphaFold 3 for this subset and computed pLDDT, epitope SASA, and inverse folding likelihood. We identified the final candidates as those in the Pareto-optimal set across all five metrics. We also examined the diversity of the selected antibody designs. We compared antibody PLM embeddings (AntiBERTy), sequence similarity (cdrDist), and structural deviation of the CDR-H3 loop (cdrRMSD) relative to the wild type (Supplement LABEL:sup:sec:computationalMetrics).
4 Results
4.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log-Likelihood
Figure 4: Spearman’s rank correlation coefficients between model log likelihood from various protein models and experimental binding affinities across multiple datasets. Models marked with * are structure-informed.
Figure 5: Proportion of top-10 ranked antibody designs achieving $\geq$5-fold affinity improvement across models and datasets. Only datasets reporting affinity as $-\!\log K_{d}$ were used. Datasets based on enrichment scores were excluded, as enrichment reflects relative sequence abundance and cannot determine fold change. Models marked with * are structure-informed.
As a result, inverse folding models achieved the highest accuracy in predicting experimental binding affinity, attaining the highest average Spearman correlation (Fig. 4) and highest 5-fold precision@10 (Fig. 5) for all inverse folding models tested. This high accuracy may stem from the model’s broader structural scope. PLMs with local structure token have consistently performed best in general protein binding tasks [41, 67], but the same did not hold for antibodies. Inverse-folding models, which encode the entire Ab–Ag complex as a single global representation, consistently outperform models that rely on local structural tokens—such as SaProt, ProSST, and ESM-3. Besides, purely autoregressive sequence models without structure like ProGen2 and ProtGPT2 likewise fail to achieve competitive accuracy on the affinity prediction task. We shuffled the chain order of autoregressive models, but it did not increase the accuracy either (Table LABEL:tab:autoregressive, Supplement LABEL:sup:sec:sec:chain_order). From a structural biology perspective, the inverse folding strategy is effective because protein function is ultimately determined by its three-dimensional structure, which is encoded by the underlying sequence. Mutations that maintain or improve structural integrity have a higher potential to enhance functional properties [63]. On the other hand, by conditioning on the full antibody-antigen structure, inverse folding models can capture long-range residue interactions and contextual features at the binding interface—both of which are critical for affinity but often missed by models relying only on local sequence information [59, 73]. Moreover, among non–inverse folding models, SaProt which leverages local structural representations achieves the second-best performance, particularly in the 5-fold precision@10 metric.
Interestingly, we also found that antibody-finetuned PLMs have varying impacts on zero-shot correlation to binding affinity. When we compare general PLM ESM-2 to its Ab-finetuned model CurrAb [5], CurrAb improved Spearman correlation by +0.074 across all datasets and precision@10 by +0.067. However, AntiFold [26], which is an ESM-IF finetuned on structure of antibodies (OAS [57]) and Ab-Ag complex (SAbDab [15]) data, deteriorated affinity correlation by -0.097 and precision@10 by -0.078 across all datasets, suggesting limited gains on Ab-specific finetuning and potential catastrophic forgetting. [^3] Note that AntiFold reports higher Spearman correlation with 1mlc dataset (0.427) – see Fig S9 [26]). This discrepancy occurs because AntiFold is evaluated using light and heavy chain variants measuring log-likelihood of the CDR regions, whereas AbBiBench excludes light chain mutants and report log-likelihood values of the entire Ab-Ag complex.
Figure 6: A. Binding potential and sequence plausibility in Phase 1. Binding potential was presented as negative binding energy changes $-\Delta\Delta G$, and $\Delta\Delta G$ is defined as $\Delta G_{variant}-\Delta G_{wild}$, where $\Delta G_{wild}=66.34$. The top 20% of variants, defined by plausibility and binding affinity, are shown in the upper right corner. B. Top 20 % variants’ CDR-H3 sequence diversity by a UMAP plot of sequence embeddings. 18 non-dominated Pareto-optimal variants, which are marked with a $\bm{\times}$ symbol, were identified by considering all five metrics on binding potential and sequence plausibility.
Figure 7: Binding potential and sequence plausibility in Phase 2. A. Sequence plausibility measured by ProteinMPNN likelihood. B. Binding potential measured by epitope SASA (presented as negative change, $-\Delta SASA$) . C. Binding potential measured by pLDDT of Ab-Ag complex structure.
4.2 Generate Antibody Variants with Strong Affinity to H1N1 Influenza Virus
We evaluated whether selected models could generate antibody variants with stronger binding affinity to H1N1 than the wild-type F045-092. Each model produced 1,500 CDR-H3 variants with up to five mutations. As DiffAb often generated wild-type sequences with altered structures, we excluded duplicates, resulting in 467 valid DiffAb variants and a total of 4,967 unique variants across all models. In Phase 1 screening, ESM-IF and SaProt produced variants with both strong binding potential and high sequence plausibility. Their average $\Delta\Delta G$ values were -29.27 and -20.79, respectively, indicating substantial improvement in binding energy (Fig. 6, LABEL:fig:phase1A). These variants also retained AntiBERTy plausibility scores (-0.663 for ESM-IF and -0.666 for SaProt) close to that of the wild type (-0.655; Fig. LABEL:fig:phase1B). On the other hands, DiffAb generated plausible sequences (-0.657) but failed to improve binding energy ($\Delta\Delta G=2.03$), suggesting mutations without affinity gain. In contrast, MEAN produced variants with improved binding ($\Delta\Delta G=-13.95$), but at the cost of lower sequence plausibility (-0.680), indicating a trade-off between biophysical fitness and evolutionary realism (Fig. LABEL:fig:phase1).
Figure 8: Relationship between structure confidence (pLDDT) and structural deviation from wild type (cdrRMSD) of CDR-H3 generated by models.
Among the top 20% of Phase 1 variants, 158, 91, and 26 candidates were selected from ESM-IF, SaProt, and MEAN, respectively. In Phase 2, SaProt achieved the highest sequence plausibility (ProteinMPNN log-likelihood = -1.406) and complex structure integrity (pLDDT = 83.22), followed by ESM-IF (-1.415, pLDDT = 82.82; Fig.  7LABEL:fig:phase2-2).
To assess the sequence diversity, we visualized CDR-H3 embeddings using AntiBERTy (Fig. 6B). MEAN variants formed a tight, isolated cluster, while ESM-IF and SaProt occupied broader, partially overlapping regions—indicating greater diversity. Structural comparison revealed that SaProt generated CDR-H3 variants closest to wild type (mean cdrRMSD = 2.28), followed by ESM-IF (2.70) and MEAN (3.25). SaProt also produced variants with higher CDR-H3 structural confidence (pLDDT), suggesting better foldability (Fig. 8LABEL:fig:diversity-cdrh3). Sequence divergence analysis using cdrDist showed SaProt variants were most similar to wild type (0.222), compared to ESM-IF (0.279) and MEAN (0.277) (Fig. 8, LABEL:fig:diversity-sw). This aligns with the average number of substitutions per CDR-H3: 2.98 for SaProt, 3.03 for MEAN, and 3.75 for ESM-IF.
5 Conclusion and Limitations
This study introduces AbBiBench, a biologically relevant and structurally informed benchmarking framework for antibody binding affinity maturation and optimization. Recognizing the limitations of traditional computational evaluation metrics, our approach explicitly incorporates antibody-antigen complex information, thereby aligning more closely with the biological realities of antibody interactions. Our results demonstrate that global structure-informed protein language model used for inverse folding methods, such as ESM-IF and ProteinMPNN, outperform other evaluated computational models, primarily due to their effective integration of structural context. In a case study focused on redesigning the F045-092 antibody for binding to the H1N1 influenza subtype, we identified 18 Pareto-optimal antibody variants with improved predicted affinity and structural integrity. These 18 variants have been successfully expressed in vitro, and wet lab experiments to quantify their binding affinity to H1N1 hemagglutinin are currently underway. The results of these assays will directly validate the computational predictions and help assess the true binding potential of the designed antibodies. Limitations of this study include the lack of currently available experimental neutralization data and the relatively small size of some benchmark datasets (Supplement LABEL:sup:sec:limitations). In future work, we plan to expand the benchmark to include additional therapeutic properties such as stability, immunogenicity, and developability, and to incorporate functional assay data (e.g., IC50) to further align model evaluations with biological outcomes.
Supplement Materials
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
IMPORTANT, please:
  • Delete this instruction block, but keep the section heading “NeurIPS Paper Checklist",
  • Keep the checklist subsection headings, questions/answers and guidelines below.
  • Do not modify the questions and only use the provided macros for your answers.
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction clearly state the contributions, inclusion of 11 antibody-antigen datasets, and the evaluation of protein models for affinity prediction and antibody design. These contributions are supported by empirical results in Sections 3 and 4.
  5. 2.
    Limitations
  6. Question: Does the paper discuss the limitations of the work performed by the authors?
  7. Answer: [Yes]
  8. Justification: A dedicated discussion of limitations is provided in Section 5 and Supplement Section LABEL:sup:sec:limitations, including the lack of experimental neutralization assays and limited dataset size for some antigens.
  9. 3.
    Theory assumptions and proofs
  10. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  11. Answer: [N/A]
  12. Justification: The paper does not present any formal theoretical results, proofs, or theorems.
  13. 4.
    Experimental result reproducibility
  14. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper?
  15. Answer: [Yes]
  16. Justification: The paper includes detailed descriptions of datasets (Section 3.1), model settings, evaluation tasks (Section 3.3), and provides full dataset/code access via Hugging Face and GitHub (Supplement Section LABEL:sup:resources).
  17. 5.
    Open access to data and code
  18. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results?
  19. Answer: [Yes]
  20. Justification: Dataset is publicly available at Hugging Face (https://huggingface.co/datasets/AbBibench/Antibody_Binding_Benchmark_Dataset) and code is available via GitHub (https://github.com/MSBMI-SAFE/AbBiBench).
  21. 6.
    Experimental setting/details
  22. Question: Does the paper specify all the training and test details?
  23. Answer: [Yes]
  24. Justification: Sections 3.23.3 and Supplement Materials Section LABEL:calculation_detailsLABEL:sup:sec:case_study include all necessary details on model selection, sampling strategies, and evaluation metrics.
  25. 7.
    Experiment statistical significance
  26. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  27. Answer: [Yes]
  28. 8.
    Experiments compute resources
  29. Question: Does the paper provide sufficient information on the computer resources?
  30. Answer: [Yes]
  31. Justification: As illustrated in Supplement Materials LABEL:sup:resources, model training is not required in this study. All inference tasks are conducted on a single NVIDIA H100 80GB GPU per model.
  32. 9.
    Code of ethics
  33. Question: Does the research conducted in the paper conform with the NeurIPS Code of Ethics?
  34. Answer: [Yes]
  35. Justification: The work follows ethical standards in open-source data sharing, does not involve human subjects or dual-use risks, and conforms with NeurIPS ethical principles.
  36. 10.
    Broader impacts
  37. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  38. Answer: [Yes]
  39. Justification: The manuscript highlights implications for therapeutic antibody development and model misuse risks are low given the biological specificity of antibody design (Section 5 and case study).
  40. 11.
    Safeguards
  41. Question: Does the paper describe safeguards for responsible release of high-risk data/models?
  42. Answer: [N/A]
  43. Justification: The dataset and models do not pose high risk for misuse, as they are focused on antibody sequence design for therapeutic research.
  44. 12.
    Licenses for existing assets
  45. Question: Are existing assets (code, data, models) properly credited and licensed?
  46. Answer: [Yes]
  47. Justification: All reused datasets and models are properly cited (e.g., FLAb, ProteinMPNN, SaProt, etc.) with sources listed in sections 2, 3.2 and Supplement Materials LABEL:sup:data_details.
  48. 13.
    New assets
  49. Question: Are new assets well documented and is documentation provided?
  50. Answer: [Yes]
  51. Justification: The AbBiBench dataset is introduced with detailed documentation on Hugging Face and Croissant format metadata (see Supplement Materials LABEL:sup:resources).
  52. 14.
    Crowdsourcing and research with human subjects
  53. Question: For crowdsourcing experiments and research with human subjects, does the paper include full instructions and compensation?
  54. Answer: [N/A]
  55. Justification: The work does not involve any human participants or crowdsourcing.
  56. 15.
    Institutional review board (IRB) approvals
  57. Question: Does the paper describe IRB approval for human subject research?
  58. Answer: [N/A]
  59. Justification: No research involving human subjects was conducted.
  60. 16.
    Declaration of LLM usage
  61. Question: Does the paper describe LLM usage if it is an original or non-standard component?
  62. Answer: [N/A]
  63. Justification: This study evaluates pre-trained protein language models (PLMs), which are specialized models trained on protein sequences, not general-purpose large language models (LLMs). No LLMs were used in developing or implementing the core methods of this research.