### v2

# AbBiBench: A Benchmark for Antibody Binding Affinity Maturation and Design

Abstract
We introduce AbBiBench (Antibody Binding Benchmarking), a benchmarking framework for antibody binding affinity maturation and design. Unlike previous strategies that evaluate antibodies in isolation, typically by comparing them to natural sequences with metrics such as amino acid recovery rate or structural RMSD, AbBiBench instead treats the antibody–antigen (Ab–Ag) complex as the fundamental unit. It evaluates an antibody design’s binding potential by measuring how well a protein model scores the full Ab–Ag complex. We first curate, standardize, and share more than 184,500 experimental measurements of antibody mutants across 14 antibodies and 9 antigens—including influenza, lysozyme, HER2, VEGF, integrin, Ang2, and SARS-CoV-2—covering both heavy-chain and light-chain mutations. Using these datasets, we systematically compare 15 protein models including masked language models, autoregressive language models, inverse folding models, diffusion-based generative models, and geometric graph models by comparing the correlation between model likelihood and experimental affinity values. Additionally, to demonstrate AbBiBench’s generative utility, we apply it to antibody F045-092 in order to introduce binding to influenza H1N1. We sample new antibody variants with the top-performing models, rank them by the structural integrity and biophysical properties of the Ab–Ag complex, and assess them with in vitro ELISA binding assays. Our findings show that structure-conditioned inverse folding models outperform others in both affinity correlation and generation tasks. Overall, AbBiBench provides a unified, biologically grounded evaluation framework to facilitate the development of more effective, function-aware antibody design models.
1 Introduction
Antibodies are critical components of the adaptive immune system, functioning primarily by recognizing and binding specifically to antigens such as pathogens or aberrant cells. This specific recognition is facilitated through complementary regions: the antibody’s paratope and the antigen’s epitope. Improving antibody–antigen affinity boosts therapeutic potency and is critical for drug discovery. Developing a therapeutic monoclonal antibody depends on multiple factors, including expression, stability, immunogenicity, aggregation, and binding affinity Chungyoun et al. (2024); Jain et al. (2017). Among these, binding affinity is the most critical determinant of therapeutic efficacy, as it directly influences antibody potency. Therefore, increasing binding affinity between antibody and antigen has been a crucial process in therapeutic antibody development.
Traditional antibody discovery methods, such as phage display technology Marks et al. (1992); McCafferty et al. (1990); Smith (1985) and animal immunization Green et al. (1994); Köhler & Milstein (1975), employ iterative cycles of mutation and selection to progressively improve binding affinity. These methods have significantly advanced therapeutic antibody discovery but face challenges due to the initially limited diversity of libraries in phage display or the narrow naive B cell repertoire available in animal immunization models, restricting the comprehensive exploration of potential high-affinity antibody variants (Fig. 1). Machine learning-based antibody design approaches complement these experimental techniques by efficiently navigating the vast search space (Fig. 1) and proposing high-affinity antibody variants that may not be readily accessible through experimental approaches alone.
Figure 1: Antibody design space. (a) Traditional in vitro screening explores a limited antibody library. (b) Protein machine learning models explore broader mutational space. $\bullet$: tested antibodies, $\bullet$: lead antibody, $\circ$: model-generated variants.
Recent advances in machine learning, especially with protein language models and structure-based generative models Dauparas et al. (2022); Hayes et al. (2025); Høie et al. (2025); Hsu et al. (2022); Kong et al. (2022; 2023); Li et al. (2024); Luo et al. (2022); Malherbe & Ucar (2024); Martinkus et al. (2023); Ruffolo et al. (2021); Shanker et al. (2024); Su et al. (2023); Watson et al. (2023); Wu & Li (2023), have shown promise in antibody design. However, common evaluation metrics like amino acid recovery rates or structural RMSD to natural antibodies do not adequately capture biological relevance in antibodies. In general protein design, comparing to the closest natural variant is a reasonable validation strategy because protein mutations are driven by strong evolutionary pressure. By contrast, antibody generation involves stochastic recombination and hypermutation, yielding extreme diversity. Even antibodies to the same antigen often show little sequence similarity unless clonally related. As such, evaluating designed antibodies by how closely they resemble naturally occurring ones overlooks the fundamental biology of antibody generation. This calls for new evaluation criteria that better reflect the functional goals of antibody engineering, rather than assumptions borrowed from general protein design.
From a structural biology perspective, binding affinity is determined not just by the antibody sequence, but by the quality of the interface it forms with the antigen. High-affinity binding typically arises from antibody-antigen (Ab-Ag) complexes that exhibit structural integrity Shanker et al. (2024) – meaning they are stable, well-packed, and maintain favorable conformations with minimal strain. Structural integrity ensures optimal shape and chemical complementarity at the binding interface. Antibodies that form such stable complexes with their targets resemble naturally occurring Ab-Ag complexes such as those collected in structural databases like SAbDab Dunbar et al. (2014). Recent machine learning models Dauparas et al. (2022); Hsu et al. (2022); Li et al. (2024); Luo et al. (2022); Malherbe & Ucar (2024); Su et al. (2023); Wu & Li (2023) have shown success in learning Ab‑Ag complex sequence-structure patterns, enabling us to gauge whether a designed Ab-Ag complex lies within the high‑probability manifold of structurally stable, high‑affinity complexes. Therefore, incorporating the antigen into evaluation provides a more biologically grounded and functionally relevant assessment of antibody design.
To address the limitations discussed above, we introduce AbBiBench (Antibody Binding Benchmarking), a biologically relevant benchmarking framework specifically designed for improving antibody binding affinity. Rather than assessing antibodies in isolation Chungyoun et al. (2024), we consider the Ab-Ag complex as a unit for evaluation. We curated standardized data from publicly available experimental binding affinity studies, compiling 184,500 mutated antibodies across nine antigen targets to evaluate protein models for binding affinity optimization. [^1] https://huggingface.co/datasets/AbBibench/Antibody_Binding_Benchmark_Dataset We also devised and publicly shared an efficient pipeline to rank newly designed antibodies based on complex structural integrity and biophysical properties. [^2] https://github.com/MSBMI-SAFE/AbBiBench AbBiBench is curated to avoid data leakage: although wild-type antibodies or antigens may appear in public datasets, no training corpus contains the mutant antibody–antigen complexes it evaluates. By providing a rigorous, biologically grounded benchmark for antibody design, AbBiBench will accelerate methods that lead to clinically and diagnostically impactful discoveries.
Figure 2: Overview of AbBiBench benchmarks. Antibody variants with experimentally determined affinity values are curated. Data modalities include amino acid sequences, wild-type antibody-antigen complexes, and affinity scores. A diverse set of baseline models includes general protein language models and specialized antibody models. All models are evaluated on two tasks: affinity prediction and antibody redesign. Five computational metrics assess newly designed antibodies from sequence plausibility, structural integrity, and binding affinity perspectives.
2 Related Work
2.1 Measuring Binding Affinity Between Antibody and Antigen
Three main approaches for measuring binding affinity exist: experimental Abdiche et al. (2008); Jönsson et al. (1991); Livingstone (1996), biophysics-based Buß et al. (2018); Chi et al. (2024); Weitzner et al. (2017), and data-driven (machine learning) methods Dauparas et al. (2022); Evans et al. (2021); Jumper et al. (2021); Lin et al. (2023). The most direct method to measure binding affinity is through the dissociation constant (Kd), which experimentally quantifies how tightly an antibody binds to its antigen. They are costly and low-throughput, thus computational models aim to approximate the binding affinity to enable large-scale screening. Biophysics-based computational models estimate Ab-Ag binding affinity by calculating interaction energies and free energy change, such as FoldX and $\Delta\Delta G$. It provides mechanistic insights into binding affinity. Though not direct affinity estimator, structural integrity (or plausibility) of Ab-Ag complex have been used as a proxy for binding affinity, measured via structure-based confidence metrics (interface pLDDT, iPTM) Evans et al. (2021); Jumper et al. (2021) or ML-based perplexity (e.g., perplexity from AntiBERTy, ProteinMPNN, or ESM-IF) Dauparas et al. (2022); Lin et al. (2023); Ruffolo et al. (2021). Despite the efforts to measure binding affinity computationally, the reliability and accuracy of these metrics remain an ongoing challenge Johnson et al. (2025).
2.2 Prior Benchmark Studies Evaluating Binding Affinity via Model Likelihood
Alternatively, some studies retrospectively evaluate machine learning models using historical experimental protein fitness data Chungyoun et al. (2024); Ucar et al. (2024); Wang et al. (2023). These works assume that higher model likelihoods correlate with higher protein fitness, suggesting the model can generate functional sequences. ProteinGym Notin et al. (2023) exemplifies this approach by benchmarking models against experimental data on enzymatic activity, binding, expression, and stability. It measures correlation between model perplexity and fitness scores. However, ProteinGym does not include antibodies. FLAb Chungyoun et al. (2024) addresses this gap by compiling experimental measurements specifically for antibodies, including binding affinity, specificity (e.g., polyreactivity), immunogenicity, and developability metrics. While comprehensive, FLAb does not consider the antigen when evaluating binding affinity. This is a critical limitation, as antibody-antigen interactions are highly specific, and accurate affinity prediction requires modeling the Ab-Ag complex. BindingGYM Lu et al. (2024) is a a related data curation effort for protein-protein interactions. While BindingGYM is a valuable resource, our work differs in scope and construction. We are specifically interested in interactions that involve an antibody’s CDR loops and an antigen and for this reason, only 6 Ab-Ag pairs out of BindgingGYM’s 30 overlap with those we have curated here.
3 AbBiBench: An Antibody Binding Benchmark
We introduce AbBiBench, a benchmark to evaluate protein models’ ability to predict and design high-affinity antibodies. We assess zero-shot correlation between model likelihoods and experimental binding affinities (Sec. 3.2) across curated datasets (Sec. 3.1), and validate generative performance by designing CDR-H3 variants that improve F045-092 binding to H1N1 influenza (Sec. 3.3.2).
3.1 Datasets
We compiled 14 datasets containing antigen sequences, antibody heavy and light chain sequences, wild-type complex structures, and experimental binding affinity measurements. (Table 1).
For antibody structure, we focus on variable regions of heavy and light chain, where affinity conferring mutations occur. We only included datasets with at least 20 mutated antibodies to ensure the statistical significance of evaluation.
In each binding affinity study, antibody libraries are constructed through phage or yeast display, introducing mutations via deep mutational scanning or at targeted positions. Some libraries were designed computationally using machine learning Li et al. (2023); Shanehsazzadeh et al. (2023) or biophysical modeling Clark et al. (2006), resulting in variant sets ranging from 67 to 65,535. For consistency, we standardized experimental affinity measurements by taking the negative log for Kd and log for enrichment, so that higher values indicate stronger binding (Supplement LABEL:supplementary:dataset, Table LABEL:tab:dataset_details). While log enrichment is an indirect measure, it reflects how well a variant is retained after antigen-specific selection and correlates with binding strength. When normalized, it provides a scalable proxy for relative binding affinity in high-throughput screens. Notably, AbBiBench is designed to avoid data leakage: while wild-type antibodies and antigens may occur in public datasets such as OAS Olsen et al. (2022a) and SAbDab Dunbar et al. (2014), the specific antibody mutant–antigen complexes curated in our benchmark are not present in any known training corpus.
ID Antibody Antigen Variants Binding score Study
4fqi_h1 CR9114
Influenza A/
New Caledonia/20/99 (H1N1)
HC: 65,094 $\log K_{d}$ Phillips et al. (2021)
4fqi_h3 CR9114
Influenza A/
Wisconsin/67/2005 (H3N2)
HC: 65,535 $\log K_{d}$ Phillips et al. (2021)
3gbn_h1 CR6261
Influenza A/
New Caledonia/20/99 (H1N1)
HC: 1,887 $\log K_{d}$ Phillips et al. (2021)
3gbn_h9 CR6261
Influenza A/
Hong Kong/1073/1999 (H9N2)
HC: 1,842 $\log K_{d}$ Phillips et al. (2021)
aayl49 AAYL49 Spike HR2 HC: 4,312 $\log K_{d}$ Engelhart et al. (2022)
aayl49_ML AAYL49_ML Spike HR2 HC: 8,953 $\log K_{d}$ Li et al. (2023)
aayl50 AAYL50 Spike HR2 LC: 11,473 $\log K_{d}$ Engelhart et al. (2022)
aayl51 AAYL51 Spike HR2 HC: 4,320 $\log K_{d}$ Engelhart et al. (2022)
aayl52 AAYL52 Spike HR2 LC: 13,324 $\log K_{d}$ Engelhart et al. (2022)
2fjg G6.31 VEGF
HC: 2,223
LC: 2,014
log enrichment Koenig et al. (2017)
1mlc D44.1 Hen-egg-white lysozyme
HC: 1,229
LC: 865
log enrichment Warszawski et al. (2020)
1n8z trastuzumab HER2 HC: 419 $\log K_{d}$ Shanehsazzadeh et al. (2023)
1mhp AQC2 Integrin-$\alpha$-1
HC: 37
LC: 25
Both: 5
$\log K_{d}$ Clark et al. (2006)
5a12_ang2 5A12 Ang2
HC: 796
LC: 104
Both: 43
log enrichment Minot & Reddy (2024)
Table 1: Overview of the 14 Ab-Ag binding-affinity assays reported in AbBiBench, showing the number of heavy-chain mutants in each study and the respective binding metric.
3.2 Protein models
Protein modeling is a fast-evolving active research area. We selected diverse pretrained protein and/or antibody models based on originality, code availability, and structure modality (Table LABEL:tab:protein_model_summary).
Masked Language Models: Masked protein language models (MLMs) predict masked residues based on context and capture correlations between sequence motifs and higher-level functional properties, enabling their application in antibody design Hie et al. (2024); Meier et al. (2021). We evaluate several representative protein MLMs. ESM-2 Lin et al. (2023) is trained on large-scale protein sequence datasets using a masked language modeling objective. AntiBERTy Ruffolo et al. (2021) is a 12-layer BERT model trained on 57 million heavy- and light-chain sequences from antibody database (OAS Olsen et al. (2022a)). Incorporating structural information into protein language models (PLMs) improves their ability to capture spatial context beyond sequence proximity. SaProt Su et al. (2023) extends ESM-2 with structure-aware tokens from Foldseek Barrio-Hernandez et al. (2023); Van Kempen et al. (2024), embedding residue identity and local structure. ProSST Li et al. (2024) uses geometric vector perceptrons (GVP) encoder Jing et al. (2020) that discretizes local atomic neighborhoods into a compact codebook, with disentangled attention over sequence, structure, and position. ESM-3 Hayes et al. (2025), an upgraded ESM-2, is a multimodal PLM that models sequence, structure, and function through discrete token tracks. Based on such general-purpose protein MLMs, several antibody-specific MLMs have been developed, including CurrAb Burbach & Briney (2025), AbLang Olsen et al. (2022b), and IgBlend Malherbe & Ucar (2024), to capture antibody-specific mutations driven by somatic recombination and hypermutation Ruffolo et al. (2021). CurrAb, a fine-tuned ESM-2, uses curriculum learning to gradually shift from unpaired to paired OAS Olsen et al. (2022a) antibody data while preserving pre-trained knowledge on general proteins.
Autoregressive Protein Language Models: Unlike MLMs, which predict masked tokens based on bidirectional context, autoregressive PLMs generate the next token using only left-to-right context. ProGen2 Nijkamp et al. (2023) is a Transformer‐decoder model that scales up to 6.4 B parameters and is trained on $\sim\!1$ billion natural protein sequences. ProtGPT2 Ferruz et al. (2022) adopts the GPT-2 architecture with 738 M parameters and is trained end-to-end on $\sim\!50$ million protein sequences. Inverse folding models are also autoregressive but based on global structure embeddings. ProGen2 and ProtGPT2 can serve as structure-agnostic comparisons for inverse folding models.
Inverse Folding Models: Inverse folding models aim to predict amino acid sequences from a given protein structure, often in an autoregressive manner. This approach has been applied to antibody mutant design by leveraging known Ab-Ag complex structures Shanker et al. (2024), enabling sequence exploration to identify mutations that preserve or enhance complex stability and binding affinity. Widely used models include ProteinMPNN Dauparas et al. (2022) and ESM-IF Hsu et al. (2022). ProteinMPNN uses a message-passing neural network Gilmer et al. (2017) to model residue interactions before autoregressive sequence generation. ESM-IF1 combines a GVP encoder for extracting backbone-invariant features with a Transformer decoder. AntiFold Høie et al. (2025), based on ESM-IF1, is fine-tuned on antibody structures from SAbDab and OAS.
Diffusion-Based Generative Models: Diffusion models approach antibody design as a denoising process, transforming Gaussian noise into a target antibody by learning the joint distribution of atomic coordinates, orientations, and residue identities. DiffAb Luo et al. (2022) conditions on an antigen–antibody framework complex and jointly diffuses CDR sequence and structure. To assess backbone flexibility, we also evaluated fixed-backbone variants DiffAb_fixbb, a common setting in protein design Anishchenko et al. (2021); Hsu et al. (2022); Ingraham et al. (2019); Luo et al. (2022); Strokach et al. (2020); Tischer et al. (2020). AbDiffuser Martinkus et al. (2023) extends to full-atom generation with physical priors for side chains. IgDiff Cutting et al. (2025) performs de novo backbone generation by sampling variable-region backbones and then fills in the sequences using AbMPNN Dreyer et al. (2023).
CDR Imputation in Geometric Representation: Geometry-aware methods view affinity maturation as filling in missing CDRs within the explicit 3-D Ab-Ag interface. MEAN Kong et al. (2022) masks CDRs on an E(3)-equivariant residue–atom graph containing both chains and epitope; two alternating message-passing blocks jointly restore CDR sequence and backbone. dyMEAN Kong et al. (2023) upgrades this to full-atom, end-to-end design: conserved-framework initialization plus a “shadow paratope” lets the network emit paratope sequence, side-chain geometry, and binding pose in one shot. We also considered fixed-backbone versions of MEAN and dyMEAN, dubbed MEAN_fixbb and dyMEAN_fixbb, respectively.
3.3 Evaluation Tasks
Our benchmark comprises (i) zero-shot affinity prediction using retrospective experimental affinity data and (ii) antibody generation by sampling from the models.
3.3.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log‑Likelihood
To measure how well a model’s zero-shot predicted log-likelihood aligns with wet-lab verified affinity, we calculated the Spearman correlation between the model likelihood and experimentally measured binding affinity. A high correlation indicates that the model assigns a higher likelihood to strong binders, suggesting that it can identify affinity-enhancing mutations in a zero-shot setting. To evaluate how effectively a model prioritizes the most promising antibodies, we also reported 5-fold precision@10 – the proportion of top 10 ranked variants that achieve at least 5-fold improvement in binding affinity compared to the wild type. The calculation of affinity fold change is detailed in Supplement LABEL:sup:sec:computationalMetrics. We harmonized the log-likelihood computation for all models under a unified setting: the input unit is the mutant-antigen complex, and the output is the likelihood of that complex. The details of likelihood computation across different types of models are provided in Supplement LABEL:calculation_details. In addition to the model’s log likelihood, we report two biophysics-based affinity score as a baseline: binding free energy ($\Delta G$) and the relative solvent-accessible surface areas ($SASA$) of epitope residues (Sec. 3.3.2, Supplement LABEL:sup:sec:computationalMetrics). Lower values of both metrics imply stronger binding. To ensure consistent directionality with model log-likelihoods, we report $\Delta G$ and SASA-SASA.
3.3.2 Generate Antibody Variants with Strong H1N1 Influenza Affinity
To assess whether protein models can generate antibody variants with improved binding to a specific antigen, we conducted a case study using F045-092 Ohshima et al. (2011), a naturally occurring antibody that targets the hemagglutinin (HA) protein of influenza H3N2. Notably, F045-092 fails to bind the H1N1 strain California2009 due to a steric barrier at the HA receptor-binding site Ekiert et al. (2012); Simmons et al. (2023). No experimentally determined structure of the F045-H1N1 complex exists in public databases, and this particular pair has not been studied in prior antibody design literature. As such, our study represents a completely novel setting, free from data leakage. The structure used in our study was computationally predicted using AlphaFold3. We used four representative models—ESM-IF, SaProt, DiffAb, and MEAN—to redesign the CDR-H3 loop of F045-092 while allowing up to five substitutions to maintain H3N2 cross-reactivity. We focus on generating mutations in the variable heavy chain (vH) due to its high diversity from V(D)J recombination and the central role of CDR-H3 in antigen binding Chen et al. (2024). The vH often acts as a unique antigen-specific signature Davies & Riechmann (1995), while the light chain often remains relatively conserved across functional antibodies Jaffe et al. (2022), making vH the most relevant region for affinity optimization. Each model generated 1,500 CDR-H3 variants from an input consisting of the masked F045-092 sequence and a predicted complex structure with the H1N1 HA1 protein (Supplement LABEL:sup:sec:case_study). Sampling strategies varied by model, including autoregressive prediction (SaProt), greedy heuristics (ESM-IF), and diffusion-based or graph-based generation (DiffAb, MEAN; see Supplement LABEL:sup:sec:sec:design_cdrh3 for details).
We evaluated the generated antibody variants from two perspectives: sequence plausibility and binding potential. Sequence plausibility was assessed using the log-likelihood from AntiBERTy Ruffolo et al. (2021), reflecting how closely mutations align with natural antibody evolution. We also computed inverse folding likelihoods using ProteinMPNN to determine whether each sequence is compatible with its backbone structure—higher scores indicate greater foldability. Binding potential was evaluated using three structure-based metrics: (1) binding free energy ($\Delta G$) to estimate Ab–Ag interaction strength, (2) epitope SASA to measure differences in the solvent-accessible antigen surface area upon binding, and (3) complex structure confidence (AlphaFold pLDDT Abramson et al. (2024)) as a proxy for interface stability (Supplement LABEL:sup:sec:computationalMetrics).
To identify the final candidates, we used a two-phased screening approach (Fig. LABEL:fig:evaluating). In Phase 1 we evaluated all 1,500 variants per model using sequence plausibility (AntiBERTy likelihood) and $\Delta G$ – metrics that do not require structure prediction – and selected top 20% antibody variants. In Phase 2, we generated full Ab-Ag complex structures with AlphaFold 3 for this subset and computed pLDDT, epitope SASA, and inverse folding likelihood. We identified the final candidates as those in the Pareto-optimal set across all five metrics. We also examined the diversity of the selected antibody designs. We compared antibody PLM embeddings (AntiBERTy), sequence similarity (cdrDist), and structural deviation of the CDR-H3 loop (cdrRMSD) relative to the wild type (Supplement LABEL:sup:sec:computationalMetrics).
4 Results
4.1 Zero-shot Prediction of Experimental Binding Affinity using Model Log‑Likelihood
As a result, inverse folding models achieved the highest accuracy in predicting experimental binding affinity, attaining the highest average Spearman correlation (Fig. 3) and highest 5-fold precision@10 (Fig. 4) for all inverse folding models tested. This high accuracy may stem from the model’s broader structural scope. PLMs with local structure token have consistently performed best in general protein binding tasks Li et al. (2024); Su et al. (2023), but the same did not hold for antibodies. Inverse-folding models, which encode the entire Ab–Ag complex as a single global representation, consistently outperform models that rely on local structural tokens—such as SaProt, ProSST, and ESM-3. Besides, purely autoregressive sequence models without structure like ProGen2 and ProtGPT2 likewise fail to achieve competitive accuracy on the affinity prediction task. We shuffled the chain order of autoregressive models, but it did not increase the accuracy either (Table LABEL:tab:autoregressive, Supplement LABEL:sup:sec:sec:chain_order). From a structural biology perspective, the inverse folding strategy is effective because protein function is ultimately determined by its three-dimensional structure, which is encoded by the underlying sequence. Mutations that maintain or improve structural integrity have a higher potential to enhance functional properties Shanker et al. (2024). On the other hand, by conditioning on the full antibody-antigen structure, inverse folding models can capture long-range residue interactions and contextual features at the binding interface—both of which are critical for affinity but often missed by models relying only on local sequence information Orlandi et al. (2020); Wang et al. (2018).
Figure 3: Spearman’s rank correlation coefficients between model log likelihood from various protein models and experimental binding affinities across multiple datasets. Models marked with * are structure-informed.
Figure 4: Proportion of top-10 ranked antibody designs achieving $\geq$5-fold affinity improvement across models and datasets. Only datasets reporting affinity as $\log K_{d}$ were used. Datasets based on enrichment scores were excluded, as enrichment reflects relative sequence abundance and cannot determine fold change. Models marked with * are structure-informed.
Moreover, among non–inverse folding models, SaProt which leverages local structural representations achieves the second-best performance, particularly in the 5-fold precision@10 metric. Interestingly, we also found that antibody-finetuned PLMs have varying impacts on zero-shot correlation to binding affinity. When we compare general PLM ESM-2 to its Ab-finetuned model CurrAb Burbach & Briney (2025), CurrAb improved Spearman correlation by +0.074 across all datasets and precision@10 by +0.067. However, AntiFold Høie et al. (2025), which is an ESM-IF finetuned on structure of antibodies (OAS Olsen et al. (2022a)) and Ab-Ag complex (SAbDab Dunbar et al. (2014)) data, deteriorated affinity correlation by -0.097 and precision@10 by -0.078 across all datasets, suggesting limited gains on Ab-specific finetuning and potential catastrophic forgetting. [^3] Note that AntiFold reports higher Spearman correlation with 1mlc dataset (0.427) – see Fig S9 Høie et al. (2025)). This discrepancy occurs because AntiFold is evaluated using light and heavy chain variants measuring log-likelihood of the CDR regions, whereas AbBiBench excludes light chain mutants and report log-likelihood values of the entire Ab-Ag complex.
Figure 5: Binding potential and sequence plausibility in Phase 2. A. Sequence plausibility measured by ProteinMPNN likelihood. B. Binding potential measured by epitope SASA (presented as negative change, $\Delta SASA$) . C. Binding potential measured by pLDDT of Ab-Ag complex structure.
4.2 Generate Antibody Variants with Strong Affinity to H1N1 Influenza Virus
We evaluated whether selected models could generate antibody variants with stronger binding affinity to H1N1 than the wild-type F045-092. Each model produced 1,500 CDR-H3 variants with up to five mutations. As DiffAb often generated wild-type sequences with altered structures, we excluded duplicates, resulting in 467 valid DiffAb variants and a total of 4,967 unique variants across all models. In Phase 1 screening, ESM-IF and SaProt produced variants with both strong binding potential and high sequence plausibility. Their average $\Delta\Delta G$ values were -29.27 and -20.79, respectively, indicating substantial improvement in binding energy (Fig. LABEL:fig:phase1-screening, LABEL:fig:phase1A). These variants also retained AntiBERTy plausibility scores (-0.663 for ESM-IF and -0.666 for SaProt) close to that of the wild type (-0.655; Fig. LABEL:fig:phase1B). On the other hands, DiffAb generated plausible sequences (-0.657) but failed to improve binding energy ($\Delta\Delta G=2.03$), suggesting mutations without affinity gain. In contrast, MEAN produced variants with improved binding ($\Delta\Delta G=-13.95$), but at the cost of lower sequence plausibility (-0.680), indicating a trade-off between biophysical fitness and evolutionary realism (Fig. LABEL:fig:phase1).
Figure 6: ELISA $\text{OD}_{450}$ signals for 21 model-designed F045-092 mutants and controls against H1 hemagglutinin. Three antibodies were used as controls, including the wild-type F045-092 (H3-specific binder), 5J8 (H1-specific binder), and FI6v3 (a broadly neutralizing antibody for both subtypes). Diamonds represent mutants with higher $\text{OD}_{450}$ values than F045-092 against H3.
Among the top 20% of Phase 1 variants, 158, 91, and 26 candidates were selected from ESM-IF, SaProt, and MEAN, respectively. In Phase 2, SaProt achieved the highest sequence plausibility (ProteinMPNN log-likelihood = -1.406) and complex structure integrity (pLDDT = 83.22), followed by ESM-IF (-1.415, pLDDT = 82.82; Fig.  5LABEL:fig:phase2-2).
To assess the sequence diversity, we visualized CDR-H3 embeddings using AntiBERTy (Fig. LABEL:fig:phase1-screeningB). MEAN variants formed a tight, isolated cluster, while ESM-IF and SaProt occupied broader, partially overlapping regions—indicating greater diversity.
Structural comparison revealed that SaProt generated CDR-H3 variants closest to wild type (mean cdrRMSD = 2.28), followed by ESM-IF (2.70) and MEAN (3.25). SaProt also produced variants with higher CDR-H3 structural confidence (pLDDT), suggesting better foldability (Fig. LABEL:fig:diversity-rmsdLABEL:fig:diversity-cdrh3). Sequence divergence analysis using cdrDist showed SaProt variants were most similar to wild type (0.222), compared to ESM-IF (0.279) and MEAN (0.277) (Fig. LABEL:fig:diversity-rmsd, LABEL:fig:diversity-sw). This aligns with the average number of substitutions per CDR-H3: 2.98 for SaProt, 3.03 for MEAN, and 3.75 for ESM-IF.
In vitro experiments.
We performed ELISA binding assays on the 21 designed mutants against hemagglutinin (HA) from A/California/2009 (H1). The Figure 6 shows that all variants produced signals above the detection threshold for H1, indicating a gain of H1 binding activity. This suggests that our in-silico evaluation pipeline enables efficient screening of affinity-enhancing mutations from multiple models, thereby reducing the time and cost associated with experimental validation. Notably, the greatest gain in affinity against the H1 subtype was achieved by the ESM-IF derived designs supporting the conclusion of our in silico evaluation.
5 Conclusion and Limitations
This study introduces AbBiBench, a biologically relevant and structurally informed benchmarking framework for antibody binding affinity maturation and optimization. Recognizing the limitations of traditional computational evaluation metrics, our approach explicitly incorporates antibody-antigen complex information, thereby aligning more closely with the biological realities of antibody interactions. Our results demonstrate that global structure-informed protein language model used for inverse folding methods, such as ESM-IF and ProteinMPNN, outperform other evaluated computational models, primarily due to their effective integration of structural context. In a case study focused on redesigning the F045-092 antibody for binding to the H1N1 influenza subtype, we identified 21 Pareto-optimal antibody variants with improved predicted affinity and structural integrity. These 21 variants have been successfully expressed in vitro, and conduct ELISA assays to quantify their binding affinity to H1N1 hemagglutinin. The results of these assays directly validate the computational predictions and help assess the true binding potential of the designed antibodies. Limitations of this study include the lack of currently available experimental neutralization data and the relatively small size of some benchmark datasets (Supplement LABEL:sup:sec:limitations). In future work, we plan to expand the benchmark to include additional therapeutic properties such as stability, immunogenicity, and developability, and to incorporate functional assay data (e.g., IC50) to further align model evaluations with biological outcomes.
Supplement Materials